# EDBench: Large-Scale Electron Density Data for Molecular Modeling

### v1

@@ 4.3 Performance on generation task
Table 7 presents the results of HGEGNN on the ED prediction task under different ED thresholds $\rho_{\tau}$. We observe that, given the molecular structure (MS), the model can accurately predict ED values, achieving low MAE and high Pearson and Spearman correlations. This indicates that the deep learning method can significantly accelerate the generation of ED while reducing the computational cost associated with DFT calculations.
Notably, the model performance improves with increasing $\rho_{\tau}$, indicating it effectively captures high-ED regions. This aligns with chemical intuition, as high-density regions often correspond to chemically significant areas such as atomic cores and bonding regions, where the spatial patterns are more structured and consistent across molecules, making them easier for the model to learn.
@@ 4.4 Quality analysis of ED outputs from the generation task
To assess the quality of the ED data generated by HGEGNN, we employ models trained with three different random seeds, as described in Section 4.3, to generate ED5-EC data with a density threshold of $\rho_{\tau}=0.2$, denoted as G#1, G#2, and G#3. Figure 5 compares the average MAE performance of different data sources using the PointVector as baseline, where red values denote relative improvements compared to DFT-based data source. We observe that G#1, G#2, and G#3 consistently outperform the DFT-based data, indicating that HGEGNN generates high-quality ED. These demonstrate the potential of using predicted ED directly to enhance the model’s understanding of MFFs. More detailed results refer to Appendix A.8.
@@ 4.5 Ablation study on thresholds and sampling points
Due to the substantial number of ED points and their direct influence on computational efficiency, it is crucial to study the effects of ED thresholds ($\rho_{\tau}$) and sampling point counts ($\xi$) on model performance. Figure 5 shows the ablation results of PointVector under varying $\rho_{\tau}$ and $\xi$. We observe that performance does not improve proportionally with decreasing $\rho_{\tau}$ or increasing $\xi$, highlighting the importance of carefully selecting these hyperparameters to strike a balance between accuracy and computational cost. See Appendix A.9 for more details.
@@ 5 Conclusion
In this work, we explore extending molecular and atomic-level modeling in machine learning force fields (MLFFs) to the electronic level by leveraging electron density (ED). We first construct EDBench, the largest-scale dataset of its kind to date, containing 3,359,472 molecular ED computed with density functional theory (DFT). To comprehensively evaluate the potential of ED-based molecular representations, we design three categories of medium-scale benchmark tasks on top of EDBench: prediction of diverse quantum chemical properties, retrieval between molecular structures (MS) and ED, and generation of ED from MS. Extensive experiments using several state-of-the-art models demonstrate the strong potential of ED as a representation for learning in the context of MLFFs. While the EDBench project has made significant progress in the scale and quality of ED data, surpassing existing datasets, there remains room to further enhancement. In future work, we plan to expand the dataset to include higher-level functionals and material-related molecules, and to develop advanced models tailored for ED, enabling EDBench to support a broader range of scientific applications in physical chemistry.
@@ Appendix A Appendix
A.1 Detailed introduction of DFT
A.1.1 Development of quantum mechanics
The fundamental equation of quantum mechanics, the Schrödinger equation (Schrödinger and E., 1926), serves as the theoretical foundation for describing many-electron systems. The Schrödinger equation can be expressed as
@@ A.1.1 Development of quantum mechanics
Among them, Ĥ is the Hamiltonian operator, $\psi$ is the many-electron wave function, and $E$ is the total energy of the system. Solving the multi-electron Schrödinger equation directly is challenging due to the complexity arising from the multivariable wave function and electron interactions. Consequently, early researchers introduced various wave function-based approximation methods to simplify the problem. For instance, the Born–Oppenheimer approximation assumes that the nuclei are stationary relative to the electrons, which simplifies the Schrödinger equation into a function of electronic variables (Combes et al., 1981). The Hartree-Fock method further simplifies the many-electron problem into a single-electron problem by assuming that each electron moves independently in the average potential field formed by the other electrons (Szabo and Ostlund, 1996). Although the Hartree-Fock method significantly simplifies the calculations, it neglects electron correlation, leading to insufficient accuracy in some cases (Baerends et al., 1973). In addition, the Hartree-Fock method scales with the number of electrons $n$ as $\mathcal{O}(n^{4})$, its computational cost remains prohibitive for large polyatomic molecules. Density Functional Theory (DFT) is more suitable for large molecules and complex systems due to its lower computational cost ($\mathcal{O}(n^{3})$) and its ability to incorporate electron correlation effects (Gaitan and Nori, 2009).
A.1.2 Composition of effective electron potential energy
@@
A.1.3 Selection of functional and basis set
In addition, to solve the equation, it is usually necessary to select the basis set, pseudopotential, and exchange correlation functional. The basic set includes plane wave method, numerical atomic orbital method, and augmented wave method. Norm-conserving pseudopotential (NCPP), ultrasoft pesudopotential (USPP) and projector augmented wave (PAW) are common pseudopotential methods. The exchange-correlation energy functional includes the Local Density Approximation (LDA) (Perdew et al., 1996), the Generalized Gradient Approximation (GGA) (Becke, 1988), and hybrid functionals (such as B3LYP) (Lee et al., 1988). In this paper, the exchange-correlation functional used is B3LYP, and the 6-31G**/+G** basis set is selected for combination. B3LYP integrates the advantages of the Hartree-Fock method and DFT. The 6-31G**/+G** basis set enhances computational accuracy by splitting the valence electron orbitals into two sets of basis functions and further incorporating diffuse functions. This combination achieves a great balance between precision and efficiency, making it more suitable.
A.2 Example of ED visualization
@@
A.3 Discussion on the Quality of the EDBench Database
@@
To ensure the reliability and scientific utility of EDBench, we adopted a systematic and well-established protocol for electronic density (ED) calculation grounded in density functional theory (DFT) (Gaitan and Nori, 2009). The entire workflow was designed to maximize both physical fidelity and computational robustness, while minimizing potential sources of error or bias.
@@
First, all ED data were generated using Psi4 1.7, a widely used and validated open-source quantum chemistry package that supports high-accuracy ab initio and DFT calculations (Turney et al., 2012; Smith et al., 2020; Isert et al., 2022; Fliszkiewicz and Sajdak, 2023). We selected the B3LYP hybrid functional, a time-tested method known for its balance between computational efficiency and accuracy across a wide variety of molecules. Currently, B3LYP has been extensively applied in the domains of synthetic chemistry (Harimoto et al., 2025), molecular dynamics (Yang et al., 2020), phytochemistry (Nkene et al., 2025), spectroscopy (Komjati et al., 2016), medicine (Isert et al., 2022; Zhang et al., 2025) and physics (Jorge et al., 2015). This choice ensures that the resulting ED data reflect physically meaningful electron distributions rather than numerical artifacts.
@@
Meanwhile, basis sets were systematically assigned based on molecular composition, employing 6-31G** for general cases and 6-31+G** for sulfur-containing molecules to capture diffuse electronic effects (Behjatmanesh-Ardakani et al., 2011). Compared with the basic 6-31G, 6-31G**/+G** provides the description of polarizable electron distribution and electron correlation effect by adding polarization function, which improves the processing ability of molecular polarization effect. This tailored approach enhances the accuracy of electron densities, particularly in chemically relevant regions such as lone pairs, π-systems, or polarizable atoms.
@@
In addition, the reference wavefunction was selected according to spin multiplicity, with restricted Hartree-Fock (RHF) applied to closed-shell systems and unrestricted Hartree-Fock (UHF) used for open-shell systems (Yamaki et al., 2004), in line with Hund’s rule. This guarantees correct treatment of α and β spin components, reducing the risk of spin contamination and ensuring consistent modeling of open-shell species. To further control data quality, we enforced strict self-consistent field (SCF) convergence criteria before ED extraction (Woods et al., 2019; Loos et al., 2017). Electron density grids were then generated using a uniform grid spacing of 0.4 Bohr and a 4.0 Bohr padding, ensuring comprehensive spatial coverage without introducing undersampling or boundary artifacts. Additionally, a density fraction threshold of 0.85 was applied to focus on the physically relevant isosurface, filtering out low-density noise while preserving chemically meaningful features.
@@
A.4 Significance of benchmark tasks

### v2

@@ 4.3 Performance on generation task
Table 7 presents the results of HGEGNN on the ED prediction task under different ED thresholds $\rho_{\tau}$. We observe that, given the molecular structure (MS), the model can accurately predict ED values, achieving low MAE and high Pearson and Spearman correlations. This indicates that the deep learning method can significantly accelerate the generation of ED while reducing the computational cost associated with DFT calculations. Notably, the model performance improves with increasing $\rho_{\tau}$, indicating it effectively captures high-ED regions. This aligns with chemical intuition, as high-density regions often correspond to chemically significant areas such as atomic cores and bonding regions, where the spatial patterns are more structured and consistent across molecules, making them easier for the model to learn. We show the element-wise and size-resolved error statistics in Appendix A.18.
@@ 4.4 Quality analysis of ED outputs from the generation task
To assess the quality of the ED data generated by HGEGNN, we employ models trained with three different random seeds, as described in Section 4.3, to generate ED5-EC data with a density threshold of $\rho_{\tau}=0.2$, denoted as G#1, G#2, and G#3. Figure 4 compares the average MAE performance of different data sources using the PointVector as baseline, where red values denote relative improvements compared to DFT-based data source. We observe that G#1, G#2, and G#3 consistently outperform the DFT-based data, indicating that HGEGNN generates high-quality ED. These demonstrate the potential of using predicted ED directly to enhance the model’s understanding of MFFs. Notaby, the superior downstream performance of model-generated EDs does not necessarily indicate higher physical fidelity; rather, it may result from their smoother, more learnable patterns that align better with the inductive biases and optimization dynamics of downstream models than DFT-based EDs. More detailed results refer to Appendix A.19.
@@ 4.5 Ablation study on thresholds and sampling points
Due to the substantial number of ED points and their direct influence on computational efficiency, it is crucial to study the effects of ED thresholds ($\rho_{\tau}$) and sampling point counts ($\xi$) on model performance. Figure 5 shows the ablation results of PointVector under varying $\rho_{\tau}$ and $\xi$. We observe that performance does not improve proportionally with decreasing $\rho_{\tau}$ or increasing $\xi$, highlighting the importance of carefully selecting these hyperparameters to strike a balance between accuracy and computational cost. See Appendix A.20 for more details. In addition, we also performed ablation experiments on the Bohr grid spacing (Appendix A.21) and the InfoNCE temperature (Appendix A.22).
@@ 5 Conclusion
In this work, we explore extending molecular and atomic-level modeling in machine learning force fields (MLFFs) to the electronic level by leveraging electron density (ED). We first construct EDBench, the largest-scale dataset of its kind to date, containing 3,359,472 molecular ED computed with density functional theory (DFT). To comprehensively evaluate the potential of ED-based molecular representations, we design three categories of medium-scale benchmark tasks on top of EDBench: prediction of diverse quantum chemical properties, retrieval between molecular structures (MS) and ED, and generation of ED from MS. Extensive experiments using several state-of-the-art models demonstrate the strong potential of ED as a representation for learning in the context of MLFFs.
Limitations and future works. Despite the significant progress made by the EDBench project in terms of the scale and quality of electron density (ED) data, there are still limitations. For instance, the choice of density functional can be further improved, and ED-specific models can be developed to represent ED from different perspectives, such as point clouds, voxels, or images. In retrieval tasks, advanced contrastive methods (e.g., MoCo or hard negatives) may further improve performance, especially when batch negatives are too easy. While the primary goal of this work is to highlight the role of ED in MLFFs and to provide a fair and reproducible benchmark, these limitations also point toward promising directions for future work. We plan to extend the dataset to include higher-level functionals and molecules relevant to materials science, and to develop advanced representation models tailored to ED, enabling EDBench to support a broader range of applications in physical and chemical sciences.
6 Acknowledgement
This work was supported by the National Natural Science Foundation of China (grant nos. U22A2037, 62425204, 62122025, 62450002, 62432011), Grants of Ningbo 2023CX050011.
@@ Appendix A Appendix
A.1 Background Speed Reading: Electron Density, Machine-Learning Force Fields, and EDBench
A.1.1 Definition of Electronic Density and Importance of ED Data at Scale
Electron density $\rho(r)$ is the quantum-mechanical probability of finding any electron in an infinitesimal volume around the point r in three-dimensional space. Because the Hohenberg-Kohn theorem proves that the ground-state energy, bond lengths, dipole moment, reactivity and essentially every other equilibrium property of a molecule are uniquely determined by $\rho(r)$, knowing the density is equivalent to knowing the molecule’s chemistry. Yet obtaining $\rho(r)$ from first-principles Density Functional Theory (DFT) demands minutes to hours even for a small molecule; With 3.4 million molecules, the total computational cost exceeded 205,000 core-hours, equivalent to approximately 23.4 years of single-core compute time. Consequently, EDBench therefore provides the first large-scale resource in which every one of 3.36 million of molecule structures is accompanied by a consistent, high-level $\rho(r)$ grid, together with standard quantum observables such as total energy, dipole moments and frontier orbital energies, enabling machine-learning models to learn directly from the electron cloud.
A.1.2 Description of Machine-Learning Force Fields
Chemists and materials scientists traditionally predict how molecules move, react or bind by solving Newton’s equations with forces obtained from expensive quantum calculations. A machine-learning force field (MLFF) replaces this quantum step with a neural network that learns to predict forces or energies directly from atomic coordinates. Current MLFFs typically represent molecules as collections of atoms linked by bonds, angles and torsions, but they ignore the underlying electron distribution; this limits their accuracy whenever subtle electronic effects such as charge transfer or catalytic activation are involved.
A.1.3 Application of EDBench in the Community
Researchers can pre-train encoders on the proposed large-scale density data and then fine-tune them on smaller downstream tasks such as property prediction, docking or reactivity estimation; Predictive models can learn the mapping from electron density to various quantum chemical properties, such as energy and orbital energies, enabling more efficient quantum chemical property predictions. Generative models can learn a direct map from structure to $\rho(r)$ to obtain DFT-quality densities in milliseconds; Cross-modal retrieval allows an unknown experimental density—obtained, for example, from X-ray or electron diffraction—to be matched automatically against the database for inverse design. In addition, $\rho(r)$ is physically meaningful, any model that operates on it may remain chemically interpretable: one can ask where the lone pairs reside or which regions are electron-rich and receive an physically grounded answer.
A.2 Detailed introduction of DFT
A.2.1 Development of quantum mechanics
The fundamental equation of quantum mechanics, the Schrödinger equation [25], serves as the theoretical foundation for describing many-electron systems. The Schrödinger equation can be expressed as
@@ A.2.1 Development of quantum mechanics
Among them, Ĥ is the Hamiltonian operator, $\psi$ is the many-electron wave function, and $E$ is the total energy of the system. Solving the multi-electron Schrödinger equation directly is challenging due to the complexity arising from the multivariable wave function and electron interactions. Consequently, early researchers introduced various wave function-based approximation methods to simplify the problem. For instance, the Born–Oppenheimer approximation assumes that the nuclei are stationary relative to the electrons, which simplifies the Schrödinger equation into a function of electronic variables [26]. The Hartree-Fock method further simplifies the many-electron problem into a single-electron problem by assuming that each electron moves independently in the average potential field formed by the other electrons [27]. Although the Hartree-Fock method significantly simplifies the calculations, it neglects electron correlation, leading to insufficient accuracy in some cases [70]. In addition, the Hartree-Fock method scales with the number of electrons $n$ as $\mathcal{O}(n^{4})$, its computational cost remains prohibitive for large polyatomic molecules. Density Functional Theory (DFT) is more suitable for large molecules and complex systems due to its lower computational cost ($\mathcal{O}(n^{3})$) and its ability to incorporate electron correlation effects [28].
A.2.2 Composition of effective electron potential energy
@@
A.2.3 Selection of functional and basis set
To solve the equation S2, it is usually necessary to select the basis set, pseudopotential, and exchange correlation functional. The basic set includes plane wave method, numerical atomic orbital method, and augmented wave method. Norm-conserving pseudopotential (NCPP), ultrasoft pesudopotential (USPP) and projector augmented wave (PAW) are common pseudopotential methods. The exchange-correlation energy functional includes the Local Density Approximation (LDA) [30], the Generalized Gradient Approximation (GGA) [31], and hybrid functionals (such as B3LYP) [32]. In this paper, the exchange-correlation functional used is B3LYP, and the 6-31G**/+G** basis set is selected for combination. B3LYP integrates the advantages of the Hartree-Fock method and DFT. The 6-31G**/+G** basis set enhances computational accuracy by splitting the valence electron orbitals into two sets of basis functions and further incorporating diffuse functions. This combination achieves a great balance between precision and efficiency, making it more suitable.
A.3 Quantum datasets overview: physical and chemical properties
In Table S1, we not only highlights the large-scale electron density annotations provided by EDBench but also visually and statistically contrasts its physical and chemical diversity with other datasets.This addition aims to offer domain experts a clearer perspective on how EDBench complements and extends the scope of current resources, emphasizing its distinct characteristics and broad coverage. We believe this enhanced visualization will be instrumental in understanding its potential applications in quantum chemistry research.
Table S1: Quantum datasets overview about physical and chemical properties.
Datasets EDBench QM9 QH9 PubChemQC QMugs OQMD ECD
Moleculars 3359472 134K 130K 85M 665k 1.2M 140K
Conformers 3359472 134K 130K 85M 2M 1.2M 140K
Elements H, C, N, O, Ti, Ar, S, Se, He, Be, F, P, Si, Ca, Ga, Zn, Ge, Mg, B, Cl, As, Br H, C, N, O, F H, C, N, O, F H, C, N, O, P, S, F, Cl, Na, K, Mg, Ca H, C, N, O, P, S, Cl, K, Ca, Br, I Inorganic crystals Inorganic crystals
Electron density ρ √(CUBE) × × × × × √(CHGCAR)
Total Energy × × ×
7 Orbital energies (HOMO, LUMO) × × ×
HOMO-LUMO gap × × ×
Hamiltonians × × × × ×
Dipole Moment × × ×
Electronic Structure × ×
Other Energies DF-RKS Final Energy, Nuclear Repulsion Energy, One-Electron Energy, Two-Electron Energy, Exchange-Correlation Energy Zero point vibrational energy, Internal energy at 0 K/298.15 K, Free energy at 298.15 K - - Exchange-correlation energy, Nuclear repulsion energy, One-electron energy, Two-electron energy Compound formation energies -
A.4 Why not reuse the calculation results of PubCheMQC?
PCQM4Mv2 used in EDBench is a quantum chemistry dataset originally curated under the PubChemQC project[59], yet PubChemQC emphasizes ground-state electronic-structure quantities such as orbital energies[71], whereas our dataset centers on electronic density $\rho$. We provide CUBE-format $\rho$ files and NO occupancies that describe electron distribution in real space—data absent from PubChemQC. This enables direct analysis of bonding, charge transfer, and intermolecular interactions, making EDBench the large-scale, high-quality electronic-density resource PubChemQC lacks. In addition, compared to 6-31G* basis set on PubChemQC, EDBench employs the 6-31G/+G basis set with the higher computational accuracy, which adds polarization functions for hydrogen atoms and includes diffuse functions. Weak interactions, such as van der Waals forces, hydrogen bonds, and $\pi-\pi$ stacking, can be effectively described. This is more suitable for systems with high chemical accuracy requirements, such as anions and weakly interacting systems. Overall, EDBench provides the electron density resources that PubchemQC lacks and improves the DFT calculation settings.
A.5 Example of ED visualization
@@
A.6 Clarification of time costs
To clarify, the molecules in the EDBench dataset are based on DFT-optimized geometries, that is, they are optimized at the DFT level. Specifically, the molecular geometries were sourced from the PCQM4Mv2 dataset [59], where they were optimized using DFT at the B3LYP/6-31G* level. Detailed information about the DFT geometry optimization process can be found in reference [71]. Since the molecular geometries are already optimized at the quantum DFT level, we perform single-point calculations on these pre-optimized structures to obtain additional quantum chemical properties, such as electron density. This approach eliminates the need for full geometry optimization, and instead, each molecule undergoes just one SCF calculation, which is consistent with the time estimated for a single SCF cycle at the B3LYP/6-31G level on a single CPU. Thus, the 205,000 core-hours reported for 3.3 million molecules corresponds to the computational cost for these single-point SCF calculations.
A.7 Discussion on the Quality of the EDBench Database
@@
To ensure the reliability and scientific utility of EDBench, we adopted a systematic and well-established protocol for electronic density (ED) calculation grounded in density functional theory (DFT) [28]. The entire workflow was designed to maximize both physical fidelity and computational robustness, while minimizing potential sources of error or bias.
@@
First, all ED data were generated using Psi4 1.7, a widely used and validated open-source quantum chemistry package that supports high-accuracy ab initio and DFT calculations [60, 61, 56, 72]. We selected the B3LYP hybrid functional, a time-tested method known for its balance between computational efficiency and accuracy across a wide variety of molecules. Currently, B3LYP has been extensively applied in the domains of synthetic chemistry [73], molecular dynamics [74], phytochemistry [75], spectroscopy [76], medicine [56, 77] and physics [78]. This choice ensures that the resulting ED data reflect physically meaningful electron distributions rather than numerical artifacts.
@@
Meanwhile, basis sets were systematically assigned based on molecular composition, employing 6-31G** for general cases and 6-31+G** for sulfur-containing molecules to capture diffuse electronic effects [79]. Compared with the basic 6-31G, 6-31G**/+G** provides the description of polarizable electron distribution and electron correlation effect by adding polarization function, which improves the processing ability of molecular polarization effect. This tailored approach enhances the accuracy of electron densities, particularly in chemically relevant regions such as lone pairs, π-systems, or polarizable atoms.
@@
In addition, the reference wavefunction was selected according to spin multiplicity, with restricted Hartree-Fock (RHF) applied to closed-shell systems and unrestricted Hartree-Fock (UHF) used for open-shell systems [80], in line with Hund’s rule. This guarantees correct treatment of α and β spin components, reducing the risk of spin contamination and ensuring consistent modeling of open-shell species. To further control data quality, we enforced strict self-consistent field (SCF) convergence criteria before ED extraction [81, 82]. Electron density grids were then generated using a uniform grid spacing of 0.4 Bohr and a 4.0 Bohr padding, ensuring comprehensive spatial coverage without introducing undersampling or boundary artifacts. Additionally, a density fraction threshold of 0.85 was applied to focus on the physically relevant isosurface, filtering out low-density noise while preserving chemically meaningful features.
@@
In addition, post-calculation validation proceeded on two additional fronts. First, to verify that the choice of functional does not introduce systematic bias, we randomly selected 10,000 structures and recomputed their electron densities with the modern SCAN and ωB97X-D functionals while keeping the basis set fixed. Experimental result demonstrates that B3LYP yields highly consistent electron density distributions compared to more modern functionals, with minimal loss in fidelity. Second, we examined the integrity of the neural-generated densities produced by HGEGNN (Section 4.4). Replacing the original DFT grids in the ED5-EC benchmark with these generated densities actually improved downstream model performance, indicating that the neural network not only preserved chemically relevant features but also introduced a beneficial regularisation effect. Reproducibility is guaranteed through complete open release. All data files, environment specifications, checkpoint and benchmarks from EDBench will be freely available.
@@
A.8 Fidelity of different functionals to the electron density
We conducted a comparative evaluation of ED fidelity across different DFT functionals. Specifically, we randomly sampled 10,000 molecules from EDBench, and recomputed their electron densities using two higher-rung, modern functionals: SCAN and ωB97X-D, while keeping the basis set consistent with B3LYP for a fair comparison. To ensure meaningful comparison, we first parsed the resulting cube files and interpolated all EDs onto a unified grid, aligning them to the grid resolution used in B3LYP calculations. We then computed voxel-wise Root Mean Square Error (RMSE) and Pearson correlation between the B3LYP densities and those obtained with SCAN and ωB97X-D.
The results are summarized in Table S2. We observe that: (i) Between B3LYP and SCAN, the RMSE is 0.00039, with a Pearson correlation of 1.0; (ii) Between B3LYP and ωB97X-D, the RMSE is 0.00022, also with a Pearson correlation of 1.0. These findings indicate that B3LYP yields highly consistent electron density distributions compared to more modern functionals, with minimal loss in fidelity. This validates our decision to use B3LYP in the construction of EDBench and further supports its scientific utility and reliability.
Table S2: Root Mean Square Error (RMSE) and Pearson Correlation coefficient between electron densities generated by different functionals.
Functional Pair RMSE ↓ Pearson ↑
B3LYP vs. SCAN 0.00039±0.00011 1.000±0.000
B3LYP vs. ωB97X-D 0.00022±0.00014 1.000±0.000
A.9 Validation of electron densities for Ti- and Zn-containing molecules
Transition-metal-containing systems such as Ti and Zn require special consideration due to their being d-block elements and having partially filled d-orbitals. To validate the reliability of the basis set used in EDBench for these elements, we conducted a comprehensive evaluation comparing our current setup (B3LYP/6-31G**) against more advanced basis sets and functionals, including ECP-based options. Concretely, we extracted all Ti- and Zn-containing molecules from EDBench and recomputed their electron densities using 20 different functional-basis set combinations. These included 5 functionals (B3LYP, SCAN, ωB97X-D, M06, and ωB97M-V) and 4 basis sets (Def2-SVP, Def2-TZVP, Def2-TZVPP, and def2-QZVP), several of which are ECP-based and widely used for transition metal systems. We then compared the newly computed electron densities with those originally used in EDBench by computing voxel-wise RMSE and Pearson correlation. Full quantitative results are provided in Tables S3 and S4. In summary:
  • The RMSE between EDBench densities and those from higher-level configurations ranged from 0.00091 (e.g., SCAN/Def2-TZVP and SCAN/Def2-TZVPP) to 0.00551 (e.g., B3LYP/Def2-SVP).
  • The Pearson correlation coefficients across all comparisons were consistently near 1.0, indicating strong agreement in overall electron density structure.
  • These findings demonstrate that the electron densities in EDBench—though generated using a moderate basis set—remain highly consistent with those produced by more advanced, ECP-based approaches, validating the practical quality and robustness of the dataset for Ti and Zn systems.
These findings demonstrate that the electron densities in EDBench—though generated using a moderate basis set—remain highly consistent with those produced by more advanced, ECP-based approaches, validating the practical quality and robustness of the dataset for Ti and Zn systems.
Table S3: Root mean square error (RMSE) between electron densities generated by different functional and basis set combinations (Functionals: B3LYP, SCAN, ωB97X-D, M06, ωB97M-V; Basis sets: def2-SVP, def2-TZVP, def2-TZVPP, def2-QZVP).
RMSE ↓ B3LYP SCAN wB97X-D M06 wB97M-V
Def2-SVP 0.00551±0.00443 0.00162±0.00057 0.00545±0.00438 0.00349±0.00374 0.00543±0.00436
Def2-TZVP 0.00475±0.00397 0.00091±0.00015 0.00483±0.00406 0.00466±0.00384 0.00472±0.00395
Def2-TZVPP 0.00473±0.00394 0.00091±0.00015 0.00466±0.00389 0.00459±0.00378 0.00469±0.00392
def2-QZVP 0.00477±0.00394 0.00096±0.00015 0.00470±0.00388 0.00460±0.00376 0.00474±0.00393
Table S4: Pearson correlation coefficient between electron densities generated by different functional and basis set combinations (Functionals: B3LYP, SCAN, ωB97X-D, M06, ωB97M-V; Basis sets: def2-SVP, def2-TZVP, def2-TZVPP, def2-QZVP).
Pearson ↑ B3LYP SCAN wB97X-D M06 wB97M-V
Def2-SVP 0.99993±0.00007 0.99999±0.00000 0.99993±0.00006 0.99997±0.00004 0.99993±0.00006
Def2-TZVP 0.99994±0.00005 1.00000±0.00000 0.99994±0.00005 0.99995±0.00005 0.99994±0.00005
Def2-TZVPP 0.99994±0.00005 1.00000±0.00000 0.99995±0.00005 0.99995±0.00005 0.99994±0.00005
def2-QZVP 0.99994±0.00005 1.00000±0.00000 0.99994±0.00005 0.99995±0.00005 0.99994±0.00005
A.10 Statement on the maximum number of atoms
That most molecules in the current release of EDBench contain fewer than 20 heavy atoms. This design choice is deliberate and aligns with the PCQM4Mv2 distribution, which focuses on small, drug-like molecules. Prioritizing this molecular size range ensures both reliable DFT convergence and high chemical relevance for tasks such as property prediction and molecular generation.
While the current dataset emphasizes smaller molecules, EDBench already demonstrates broad elemental diversity by covering 22 distinct elements—significantly more than many existing datasets, as shown in Table 1. This highlights its potential for advancing modeling tasks beyond what traditional small-molecule benchmarks support.
Looking ahead, we plan to extend EDBench to include larger and more complex molecules, particularly those relevant to materials science and physical chemistry. To support this expansion, we will incorporate more advanced DFT functionals and element-specific basis sets to maintain accuracy and computational feasibility. We believe these enhancements will broaden the scope and applicability of EDBench across diverse scientific domains.
A.11 Significance of benchmark tasks